OpenStudy (anonymous):

Find the focus directrix and axis of the parabola with equation x^2=16y

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

the standard form is x^2 = 4py if p >0, p is focus and -p is directrix yours is x^2 = 4*4 *y , therefore p=4>0 --> focus is (0,4) and directrix is y = -p = -4

OpenStudy (anonymous):

I got the same answer, but my teacher says that there are two more which doesn't make sense..

OpenStudy (loser66):

what do you mean by 2 more?

OpenStudy (anonymous):

So the question for homework was worth 15 points, I only got 5/15 for that question, I showed all my work so I'm sure that's not it,

OpenStudy (loser66):

or your prof wants you have something like y = 1/16 x^2

OpenStudy (loser66):

another one is \[x =\pm4\sqrt{y}\]

OpenStudy (loser66):

However, they don't make sense to me, since the question ask about focus and directrix which evolve to the form x^2 = 4py . why don't you ask him?

OpenStudy (anonymous):

okay, and thank you for your help.

OpenStudy (loser66):

yw